# On thermal transpiration and thermomolecular pressure difference

### v1

@@ 2.4. Estimate of $Q_{+}$ and $Q_{-}$
$\|\phi_{\langle f,\Phi_{0}\rangle_{v}}\|_{H^{2}(dx)}\left(\lVert(I-P)f\rVert_{L^{2}(\partial\Omega^{+}_{2})}+\lVert f\rVert_{L^{2}(\partial\Omega^{+}_{1})}+\lVert\phi_{2}\rVert_{L^{2}(\partial\Omega^{-}_{2})}+\lVert f_{in}\rVert_{L^{2}(\partial\Omega^{-}_{1})}\right)$
The second term (8.13), for $k\neq i$ vanishes due to the oddness. Hence we only have the $k=i$ contribution:
Using $\Delta_{x}\phi_{\langle f,\Phi_{0}\rangle_{v}}={\langle f,\Phi_{0}\rangle_{v}}$ and $\lVert\phi_{\langle f,\Phi_{0}\rangle_{v}}\rVert_{H^{2}(dx)}\leq C\lVert\langle f,\Phi_{0}\rangle_{v}\rVert_{L^{2}(dx)}$ we obtain
We therefore deduce the lemma.

### v2

@@ Acknowledgments
$G(s)=g(x,v)+\frac{1}{\kappa}\int_{s}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt,$
and
$\left\{\begin{aligned} \frac{d}{ds}G(s)&=-\frac{1}{\kappa}Q(F,F)(x-sv)\\ &=-\frac{1}{\kappa}Q_{+}(F,F)(x-sv)+\frac{1}{\kappa}\nu_{F}(x-sv)F(x-sv)\\ &=-\frac{1}{\kappa}Q_{+}(F,F)(x-sv)+\frac{1}{\kappa}\nu_{F}(x-sv)G(s),&&\text{ in }s\in[-\tau_{-}(x,-v),\tau_{-}(x,v)].\\ G(0)&=F(x,v).\end{aligned}\right$. (172)
Note that the solution of (172) is unique and $G$ has two representations. One of them is
$G(s)=g(x,v)+\frac{1}{\kappa}\int_{s}^{\tau_{-}(x,v)}Q(F,F)(x-tv)dt.$
The other is
By uniqueness, the proof is complete. ∎
Now, we consider the following boundary value problem.
[…]
We pay attention on the following question
The mild solution of (173) is
$F(x,v)=F_{in}(q(x,v),v)+\int_{0}^{\tau_{-}(x,v)}G(x-sv,v)ds.$
Theorem 8.2.
If $F$ is a mild solution of (173) and $F,G\in L^{p}(\Omega\times\mathbb{R}^{3})$ for some $p\geq 1$, then
for any $\varphi\in C^{\infty}_{c}(\Omega\times\mathbb{R}^{3})$.
Proof.
Notice that
$=\int_{0}^{h}G(x+sv,v)ds$
Moreover, we have
Hence the proof is complete. ∎